I already asked this question, but maybe I wasn't clear enough.
Can you help me solve the equation for x?
2 11x+17-x =46
(11x+17-x is in the absolute value)

Answer by stanbon(75887)

2 11x+17-x =46
(11x+17-x is in the absolute value)
---------------
2|11x+17-x| = 46
----
|11x+17-x| = 23
---
|10x+17| = 23
----
10x+17 = 23 or 10x+17 = -23
x = 6/10 or x = 4
---
x = 3/5 or x = 4
======================
Cheers,
Stan H.